How do you solve the equation by completing the square of x^2+4x=10
Answer by algebrahouse.com(1659)

x² + 4x = 10
x² + 4x + 4 = 10 + 4 {took half of the b term, squared it, and added to both sides}
x² + 4x + 4 = 14 {combined like terms}
(x + 2)² = 14 {factored the left side in to squared form}
x + 2 = ±√14 {took the square root of each side}
x = -2 ± √14 {subtracted 2 from each side}

For more help from me, visit: www.algebrahouse.com